OpenStudy (anonymous):

What is the factored form of the expression? s^4 – 16 A. (s - 2)^2(s + 2)^2 B. (s - 2)(s + 2) C. (s - i)(s + i)(s - 2)(s + 2) D. (s - 2i)(s + 2i)(s - 2)(s + 2) A, B, C, or D? Please explain (:

OpenStudy (anonymous):

s^2 - 16 = s^2 - 4^2 again the same formula a^2 -b^2 (s-4) (s+4)

OpenStudy (anonymous):

because s-4 is again (s+2)(s-2)

OpenStudy (anonymous):

\[\huge a^2 - b^2 = (a +b)(a-b)\]

OpenStudy (anonymous):

sorry answer is "D" (s - 2i)(s + 2i) =(s+4)

OpenStudy (anonymous):

\(\large s^4-16=(s^2)^2-4^2=(s^2-4)(s^2+4) \) that first factor is factorable using difference of squares. that second one is factorable involving "sum of squares"

OpenStudy (anonymous):

Thanks a lot guys (: